OpenStudy (anonymous):

If the legs of a right triangle are 8 and 10, find the length of the hypotenuse. (it has to be put into radical form) something times the square root of something else.

OpenStudy (anonymous):

\[\sqrt{8^2+10^2}=2 \sqrt{41} \]